OpenStudy (anonymous):

6. f(x) = and h(x) = x-2 . f(-3) = ? h(-3) = ? limx→-3f(x) = ___ limx→-3h(x) = ___

OpenStudy (anonymous):

6. f(x) = (x^2 + x - 6)/ (x+3)

OpenStudy (lalaly):

\[\lim_{x \rightarrow -3}{x-2} = -3-2=-5\]

OpenStudy (lalaly):

\[\lim_{x \rightarrow -3}{\frac{x^2+x-6}{x-3}}=\]\[\lim_{x \rightarrow -3}{\frac{(x+3)(x-2)}{x+3}}\]now cancel out the x+3 \[\lim_{x \rightarrow -3}{x-2} = -3-2 = -5\]